Compute performance eating of Juan in a given task with the Following information.

*Normal Time= 5 hours

*Standard Personal Time = 30 minutes

*Standard Allowance for Delay=10 minutes

*Actual time of work= 450 minutes


Compute the efficiency and utilization rate of a plant with the given data


1
Expert's answer
2021-05-17T12:18:02-0400

Efficiency = "\frac{standard or normal labor hours}{Actual time for work} \times 100"

= standard normal labor hours "\div" Actual time for work "\times" 100


standard normal labor hours( 5"\times 60" )= 300 minutes


Actual time for work = 450 minutes


Efficiency = "\frac{300}{450} \times 100= 66.67" %


Utilization Rate = "\frac{ billabe time }{total hours availabe} \times 100"


Billable time = Standard time normal time - standard personal time- delays in time


= 300- 30 -10 = 260 minutes


Utilization Rate = "\frac{ 260 }{300} \times 100 = 86.67" %